OpenStudy (anonymous):

log7 + log (n - 2) = log 6n

OpenStudy (mary.rojas):

If 7 is the base, log(7) = 1 But the calculation is the same, whatever the base is... log(7) = log(6n) - log(n-2) = log[6n / (n-2)] As the application log is a bijection : 7 = 6n / (n-2) 7n - 14 = 6n n = 14 Check : log 7 + log 12 = log 84 OK as 7 x 12 = 84 Hope it helped...

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif I'm assuming you're using base10, thus look at rule 1 on that picture after that you'd use rule 2

OpenStudy (jdoe0001):

\(\bf log(7) + log (n - 2) = log(6n)\\ \color{blue}{\text{using rule 1 on that picture}}\\ log(7) + log (n-2) \implies log(7(n-2))\\ log(7(n-2))= log(6n)\\ \color{blue}{\text{substract log(6n) from both sides}}\\ \color{blue}{\text{then use rule 2 on that picture}} \)

OpenStudy (anonymous):

I don't understand this at all.